| <p><strong>08:27</strong></p>
<p>No run going.</p>
<p>Temperatures OK - attach</p>
<p> </p>
<p>Attempting to bias channel 3 again. Slowly ramping up.</p>
<p>V = 60, I = 5.3</p>
<p>V = 70, I = 6.1</p>
<p>V = 80, I = 9.0 --> 6.2 (unstable). Decreasing bias.</p>
<p>V = 70, I = 6.0 (stable again)</p>
<p>Leaving at 70 V for a while to condition (?)</p>
<p> </p>
<p>Resetting the gnuplot graph. Old caen_current.dat moved to G22-205-caen_current.dat and copied in /TapeServer/G22-205</p>
<p> </p>
<p><strong>10:10</strong></p>
<p>Temperatures OK</p>
<p>Leakage current on channel 3 showed some issues, back to stable. See attached.</p>
<p> </p>
<p><strong>11:00</strong></p>
<p>Temps OK</p>
<p>Leakage current stable.</p>
<p>Increased bias on channel 3 to 80 V. I = 6.2 uA stable. Leaving for a while.</p>
<p> </p>
<p><strong>12:45</strong></p>
<p>Temps OK</p>
<p>Leakage current stable</p>
<p>Increased bias on channel 3 to 90 V. I = 6.3 stable. Leaving for a while.</p>
<p> </p>
<p><strong>13:30</strong></p>
<p>Temps OK</p>
<p>Leakage current stable.</p>
<p>Increased bias on channel 3 to 100V. I = 6.3 stable. Hopefully it holds until the end of the beam time.</p>
<p> </p>
<p><strong>20:00</strong></p>
<p>Leakage current on channel 3 had issues again and bias was reduced to 50 V. Remained stable. Turned to 60 V now. Still stable. See attachement.</p>
<p>FEEs on detector for -> MEC</p>
<p>Temps OK - attach</p>
<p>Checks as before - attach</p>
<p>Rates OK (four fees not counting as expected) - attach</p>
<p> </p>
<p>Deuterium beam in the ring at 2 MeV/A approx 2E8 pps, significant H2+ contamination.</p>
<p>Deuterium target working, fluctuating around 5E12-1E13 atoms/cm2</p>
<p>Vacuum in YR09 - 3E-10 mbar</p>
<p> </p>
<p>Beginning beam-target overlap measurements soon.</p> |